The rate of burning of fuel in a rocket is 50g/s and comes out with velocity 4000 m/s.The force exerted by gas on rocket will be


a.200 N


b.250 N


c.2500000 N


d.25000 N

Grade:12th Pass

1 Answers

KARAN ACHARYA
12 years ago

FORCE EXERTED = CHANGE IN MOMENTUM / TIME

rate of burning fuel = 50g/s = 50/1000 kg/s

so F= (50 X 4000) / 1000 = 200 

 

Ans a.200N